OpenStudy (anonymous):

Help: 1.5m^6(-2m^2)4 !

OpenStudy (anonymous):

Hints : ================== \[\Huge (a\times c)^x = a^x\times c^x\] \[\Huge x^a\times x^c=x^{a+c}\]

OpenStudy (anonymous):

Can you solve it now ?

OpenStudy (anonymous):

Noo.. Not Good With These Problems

OpenStudy (anonymous):

Hmm, i assume this is your problem : \[\Large 1.5m^6(-2m^2)^4 \] Right ?

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

Hmmmm, First Distribute the power \[\Large (-2m^2)^4\] =\[\Large (-2^2 \times m^{2 \times 4})\]= \[\Large -4 \times m^8 \] Then Combine like terms : \[\Large \color{Red}{1.5m^6 \times -4m^8} \] = \[\Large \color{Blue}{-6m^{6+8}=-6m^{14}}\]

OpenStudy (anonymous):

\[\Huge Got~it~!?\]

OpenStudy (anonymous):

Yess

OpenStudy (anonymous):

good